A weather balloon with a volume of 211 L is launched at 20.0°C at sea level, where the atmospheric pressure is 1.00 atm. The balloon rises to an altitude of 2.00 x 104m, where atmospheric pressure is 61.0 mmHg and the temperature is 210.0 K. What is the volume of the balloon at 2.00 x 104m? 
Answer #1

P1 = 1.00 atm , T1 = 20.0°C = 293.15 K , V1 = 211 L

P2 = 61.0 mmHg or 0.0803 atm , T2 = 210.0 K , V2

Use

P1V1/T1 = P2V2/T2

= (1.00*211)/293.15 = 0.0803*V2 / 210.0

= V2 = 1882.33 L

If you have any questions please comment

If you satisfied with the solution please rate it thanks